OpenStudy (anonymous):

Hi. i'm trying to find the exact value of the infinite sum of n(2/3)^n from 1 to infinity. Is this a geometric series?

jhonyy9 (jhonyy9):

the first one will be 1*(2/3)^1 = ?

OpenStudy (thomas5267):

Is this the sum you are looking for? \[ \sum_{n=1}^\infty n\left(\frac{2}{3}\right)^n \]